# Estimating treatment effects with competing intercurrent events in randomized controlled trials

### v1

@@ 5 Simulation
Next, we evaluate two ad-hoc methods commonly used in clinical trials with ICEs, demonstrating through simulations that they are biased in estimating the clinically relevant causal parameter. The first method is the non-responder imputation method that assigns an outcome as 00 whenever an ICE occurs. It is equivalent to using the composite outcome strategy that treats all ICEs as treatment-related, leading to a composite outcome equal to 00. The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates. We generate data under the “$\mu\_S$_wrong regime, ensuring correct specification of both the propensity score $e(X)$ and the treatment-unrelated ICE survival model $G_{a}(t\mid X)$ for $a=0,1$. Based on $1000$ Monte Carlo simulations with a sample size of $n=1000$, the inverse probability-weighting estimator using the non-responder imputation method exhibits a finite sample bias of $-1.195$ with a standard deviation of $0.077$, and that using the hypothetical strategy for all ICEs has a bias of $1.188$ with a standard deviation of $0.388$. Both ad-hoc methods yield inconsistent estimators.
6 Application
In this section, we re-analyze data from two immunology trials to study the effect of baricitinib on Systemic Lupus Erythematosus (Morand et al.,, 2023; Petri et al.,, 2023). Both trials are double-blinded, randomized, placebo-controlled phase-3 trials. There are three treatment arms: 2mg baritinib, 4mg baritinib, and placebo. The primary endpoint of interest in both studies is the Systemic Lupus Erythematosus Responder Index 4 (SRI4) measured at week 52. It is a binary composite responder index based on both the level of improvement in disease activity and an indicator of no worsening of the overall condition or the development of substantial disease activity in new organ systems. The baseline covariates used in model-fitting include the region each patient is from, the baseline corticosteroid usage, and the baseline value of the Physician Global Assessment score measured at week 0.
There are significant numbers of observations with ICEs happening in both trials (218/760 and 211/775). We classify the observed intercurrent events into two types: those likely to be correlated with the potential outcomes such as adverse events (82.6% and 84.4%), and those conditionally independent of the potential outcomes given covariates and the treatment arm such as withdrawal due to relocation (17.4% and 15.6%). We compare both 2mg and 4mg treatment arms with the placebo group separately. We implemented all four estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, $\hat{\tau}^{\textup{aug}}$, and $\hat{\tau}^{\textup{eif}}$. We used logistic regression for estimating the propensity score and outcome models and used Cox proportional hazards regression to estimate the survival function and censoring model. We used the nonparametric bootstrap with 500 iterations to compute the variance estimator and conduct inference. We summarize the point estimators, their estimated standard errors, and the $p$-values in Table 4.
Table 4: Treatment effect of Baricitinib on the primary endpoint SRI4. The first four columns report the results of Trial 1 and the last four columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the composite outcome average treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on all four estimators.
Trial 1 (Petri et al.,, 2023) Trial 2 (Morand et al.,, 2023)
2mg point 0.030 0.029 0.026 0.026 0.019 0.019 0.022 0.022
se 0.042 0.043 0.042 0.043 0.042 0.043 0.043 0.042
$p$-value 0.479 0.504 0.534 0.540 0.643 0.662 0.602 0.606
4mg point 0.113 0.120 0.115 0.113 $-0.002$ $-0.002$ $-0.002$ $-0.002$
se 0.046 0.046 0.046 0.046 0.042 0.042 0.042 0.042
$p$-value 0.013 0.008 0.012 0.013 0.961 0.962 0.966 0.969
In both trials and both treatment arms, the point estimators are close to each other in most of the cases, and the statistical inference based on different estimators is remarkably robust across methods. Our results suggest a significantly positive effect of 4mg baricitinib treatment on SRI4 in trial 1, although that in trial 2 is insignificant with a slightly negative point estimator. Neither trial observes a significant effect of 2mg baricitinib treatment on SRI4.
Next, we use the two commonly used ad-hoc methods discussed at the end of Section 5 and compare them with our proposed methods. Table 5 reports the estimated treatment effect of both treatment arms across the two trials using the inverse probability-weighting estimator. The non-responder imputation method yields smaller estimates than our proposed methods most of the time, although the difference is modest due to the relatively low proportion of treatment-unrelated ICEs in both trials. Nonetheless, our proposed method targets a clinically more meaningful causal parameter. The second ad-hoc method naively applies the hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on the observed covariates. Such independence assumption is untestable and likely to be violated in clinical trials. For instance, patients discontinued treatment due to adverse events or lack of efficacy. As shown in Table 5, this method produces estimated treatment effects substantially different from the obtained results using our proposed methods, leading to misleading clinical conclusions. It aligns with the large sample bias it exhibits in the simulation study.
Table 5: Treatment effect of Baricitinib on the primary endpoint SRI4 using non-responder imputation (NRI) and hypothetical strategy (HS). The first two columns report the results of Trial 1 and the last two columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the estimated treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on the inverse probability-weighting estimator.
Trial 1 Trial 2
NRI HS NRI HS
2mg point 0.039 0.011 0.010 0.022
se 0.043 0.049 0.042 0.050
$p$-value 0.362 0.815 0.817 0.654
4mg point 0.093 0.073 $-0.008$ 0.044
se 0.044 0.050 0.041 0.051
$p$-value 0.036 0.147 0.847 0.382
7 Discussion
In this study, we address the challenges posed by ICEs in RCTs by proposing methods to handle competing ICEs. We classify ICEs into treatment-related and treatment-unrelated events and apply different strategies for the estimation of a clinically meaningful causal effect. For treatment-related ICEs, we use the composite outcome approach, assigning values that capture treatment failure. For treatment-unrelated ICEs, we adopt a hypothetical strategy that assumes their occurrence is conditionally independent of the potential outcome given covariates. Our key insight is that patients discontinue treatment following the first ICE that occurs, which in turn censors the potential occurrence of ICEs of another type. In this paper, we rigorously address the issue of competing ICEs by employing a formal causal inference framework. We provide identification formulas for the causal parameter of interest, derive its semiparametric efficient influence function, and propose multiple estimators. While our proposed framework provides a rigorous and flexible approach to handling competing ICEs, several challenges and extensions remain open for future research.
7.1 Data collection and ICE classification
Our proposed methods have broad applicability across various therapeutic areas, including immunology, oncology, and cardiology, where treatment discontinuation and other ICEs frequently occur. However, the effectiveness of these approaches relies on the accurate classification of ICEs, which should be performed in collaboration with clinicians to ensure clinical relevance in trial analyses. In our application studies, the two types of ICEs were classified based on a manual review of the detailed comments collected at the clinic. A modernized case report form for a more accurate collection of reasons for treatment discontinuation is required for each application of the proposed method in clinical trials. Fortunately, a cross-industry PHUSE working group is tackling this problem, and a recommended new case report form will be available soon (PHUSE Working Group,, 2024).
7.2 Random $K$
In practice, the measurement of the endpoint of interest $Y$ does not necessarily happen at a fixed point $k$. For instance, the time $K$ that a patient visits the clinic and takes the measurement can be treated as a random variable independent of treatment $A$, potential outcomes $Y(a)$, and both types of ICEs $T(a)$ and $C(a)$ for $a=0,1$. For a random $K$, the analogous formulas to both the previous identification formulas (2) and (3) hold, with a replacement of $k$ to $K$, i.e.,
If $K$ is a pre-treatment covariate that we observe, the previous analyses carry over. If $K$ is also treated as a post-treatment variable, we can construct a weighting estimator following (10). However, (9) ceases to serve as an identification formula unless $S_{a}(K\mid X)$ is identified for $a=0,1$. Considering the presence of three competing time-to-event random variables, it becomes necessary to employ competing risks models. We defer this analysis to future work.
7.3 Comparison between composite variable and principal stratification
We propose to use a combination of the composite variable and hypothetical strategies to deal with ICEs in this project. Following ICH E9 , 2019 (R1), we can also use the combination of principal stratification and hypothetical strategies to deal with the treatment-related and treatment-unrelated ICEs, respectively. Using the principal stratification strategy will target different causal parameters and rely on additional identification assumptions such as monotonicity and principal ignorability of the two potential values of the AE time (Ding and Lu,, 2017; Jiang and Ding,, 2021; Jiang et al.,, 2022). We leave this for future work.
Supplementary Material
We provide proof of all theorems and propositions in the supplementary material.
Appendix A Proofs
A.1 Proof of Theorem 1
@@
First, equation (2) holds because
LHS $=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]$
@@
$=$ $E\{E(Y\mid T>k,C>k,X,A=1)\textup{pr}(T>k\mid X,A=1\}=\textup{RHS},$
where the first equality is by the law of iterated expectations, the second equality is by the randomization assumption 1, and the fourth equality is by the censoring at random assumption 2.
Next, equation (3) holds because
RHS $=$ $E\left[E\left\{\frac{AY1(T\wedge C>k)}{e(X)\textup{pr}(C>k\mid X,A=1)}\mid X\right\}\right]$
@@
$=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]=\textup{LHS},$
where the first equality is by the law of iterated expectations, the third equality is by the randomization assumption 1, the fourth equality is by the censoring at random assumption 2, and the last equality is again by the law of iterated expectations. $\square$
A.2 Proof of Proposition 1
We prove that when the censoring model is correctly specified, $\hat{\tau}^{\textup{aug}}$ is doubly robust in the sense that it is consistent to $\tau$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. Want to show
if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$ holds.
$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$
@@ A.2 Proof of Proposition 1
where the first equality is by the law of integrated expectations and the second equality follows from similar derivations as in the proof of Theorem 1. The last line is equal to $\mu_{1}$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. $\square$
A.3 Proof of Theorem 2
@@ A.4 Proof of Theorem 3
$\mu_{1}=E\left[\frac{A}{e^{*}(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}+\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\right\}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$ (S2)
if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
Denote
@@ A.5 Proof of Theorem 4
Next, we show that Assumption 5 is a sufficient condition for $R_{11}=o_{P}(n^{-1/2})$ and $R_{12}=o_{P}(n^{-1/2})$. In the following discussion, we suppress the dependency of the nuisance functions and their estimated values on $X$ for ease of notation. By the facts that $e\hat{\mu}_{1}\hat{S}_{1}(k)/\hat{e}$ is bounded and $1/\{\hat{S}_{1}(t)\hat{G}_{1}(t)\}$ is bounded for any $t\leq k$, we employ the Cauchy-Schwarz inequality to upper bound the term $R_{11}$ by a constant times $\int_{0}^{k}||S_{1}(t)-\hat{S}_{1}(t)||_{2}||G_{1}(t)-\hat{G}_{1}(t)||_{2}\textnormal{d}t$, where $||\cdot||_{2}$ denotes the $L_{2}(P)$ norm. It follows that Assumption 5 guarantees $R_{11}=o_{P}(n^{-1/2})$. For $R_{12}$, we further suppress the dependency of $S_{1}(k\mid X)$, $G_{1}(k\mid X)$ and their estimators on both $k$ and $X$, and rewrite the term as $R_{12}=E[\{\hat{e}(G_{1}-\hat{G}_{1})+G_{1}(e-\hat{e})\}\{\hat{\mu}_{1}(S_{1}-\hat{S}_{1})+S_{1}(\mu_{1}-\hat{\mu}_{1})\}/(\hat{e}\hat{G}_{1})]$. Because $\hat{e}/\hat{G}_{1}$ is bounded, by the Cauchy-Schwarz inequality, the first term in $R_{12}$ is upper bounded by $||G_{1}-\hat{G}_{1}||_{2}||S_{1}-\hat{S}_{1}||_{2}=o_{P}(n^{-1/2})$ by Assumption 5. The other three terms in $R_{12}$ can be similarly bounded. Thus, $R_{1}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$.
Symmetric arguments imply the analogous term for the control arm satisfies $R_{0}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$. Therefore, results in Theorem 4 follow from the central limit theorem. $\square$

### v2

@@ B.2 Simulation results
Next, we evaluate two ad-hoc methods commonly used in clinical trials with ICEs, demonstrating through simulations that they are biased in estimating the clinically relevant causal parameter. The first method is the non-responder imputation method that assigns an outcome as 00 whenever an ICE occurs. It is equivalent to using the composite outcome strategy that treats all ICEs as treatment-related, leading to a composite outcome equal to 00. The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates. We generate data under the “$\mu\_S$_wrong” regime, ensuring correct specification of both the propensity score $e(X)$ and the treatment-unrelated ICE survival model $G_{a}(t\mid X)$ for $a=0,1$. Based on $1000$ Monte Carlo simulations with a sample size of $n=1000$, the inverse probability-weighting estimator using the non-responder imputation method exhibits a finite sample bias of $-1.195$ with a standard deviation of $0.077$, and that using the hypothetical strategy for all ICEs has a bias of $1.188$ with a standard deviation of $0.388$. Both ad-hoc methods yield inconsistent estimators.
Appendix C Proofs
C.1 Proof of Theorem 1
@@
First, equation (2) holds because its left-hand side equals
$E[E\{Y(1)1(T(1)>k)\mid X\}]$
@@
$=$ $E\{E(Y\mid T>k,C>k,X,A=1)\textup{pr}(T>k\mid X,A=1\},$
which is equal to the right-hand side of (2), where the first equality is by the law of iterated expectations, the second equality is by the randomization assumption 1, and the fourth equality is by the censoring at random assumption 2.
Next, equation (3) holds because its right-hand side equals
$E\left[E\left\{\frac{AY1(T\wedge C>k)}{e(X)\textup{pr}(C>k\mid X,A=1)}\mid X\right\}\right]$
@@
$=$ $E[E\{Y(1)1(T(1)>k)\mid X\}],$
which is equal to the left-hand side of (3), where the first equality is by the law of iterated expectations, the third equality is by the randomization assumption 1, the fourth equality is by the censoring at random assumption 2, and the last equality is again by the law of iterated expectations. $\square$
C.2 Proof of Proposition 1
We prove that under the correct specification of the censoring model, $\hat{\tau}^{\textup{aug}}$ is doubly robust in the sense that it is consistent for $\tau$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. We have
$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$ (S1)
@@ C.2 Proof of Proposition 1
where the first equality follows from the law of integrated expectations and the second equality follows from similar derivations as in the proof of Theorem 1. Now observe that the final expression of (S1) is equal to $\mu_{1}$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. Therefore, we have
if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$ holds. A parallel argument applies to the control counterpart. Therefore, combining the treated and control components, we conclude that $\hat{\tau}^{\textup{aug}}$ is consistent for $\tau$ if either the propensity score model is correct, or the outcome and survival models for the treatment-related ICE are both correct. This establishes the double robustness of $\hat{\tau}^{\textup{aug}}$. $\square$
C.3 Proof of Theorem 2
@@ C.4 Proof of Theorem 3
$\mu_{1}$ $=$ $E\left[\frac{A}{e^{*}(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}+\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\right\}\right.$ (S3)
$\quad\ \left.-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$
if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$. Define
@@ C.5 Proof of Theorem 4
Next, we show that Assumption 5 is a sufficient condition for $R_{11}=o_{P}(n^{-1/2})$ and $R_{12}=o_{P}(n^{-1/2})$. In the following discussion, we suppress the dependency of the nuisance functions and their estimated values on $X$ for ease of notation. By the facts that $e\hat{\mu}_{1}\hat{S}_{1}(k)/\hat{e}$ is bounded and $1/\{\hat{S}_{1}(t)\hat{G}_{1}(t)\}$ is bounded for any $t\leq k$, we employ the Cauchy-Schwarz inequality to upper bound the term $R_{11}$ by a constant times
$\int_{0}^{k}||S_{1}(t)-\hat{S}_{1}(t)||_{2}||G_{1}(t)-\hat{G}_{1}(t)||_{2}\textnormal{d}t,$
where $||\cdot||_{2}$ denotes the $L_{2}(P)$ norm. It follows that Assumption 5 guarantees $R_{11}=o_{P}(n^{-1/2})$. For $R_{12}$, we further suppress the dependency of $S_{1}(k\mid X)$, $G_{1}(k\mid X)$ and their estimators on both $k$ and $X$, and rewrite the term as
$R_{12}=E[\{\hat{e}(G_{1}-\hat{G}_{1})+G_{1}(e-\hat{e})\}\{\hat{\mu}_{1}(S_{1}-\hat{S}_{1})+S_{1}(\mu_{1}-\hat{\mu}_{1})\}/(\hat{e}\hat{G}_{1})].$
Because $\hat{e}/\hat{G}_{1}$ is bounded, by the Cauchy-Schwarz inequality, the first term in $R_{12}$ is upper bounded by $||G_{1}-\hat{G}_{1}||_{2}||S_{1}-\hat{S}_{1}||_{2}=o_{P}(n^{-1/2})$ by Assumption 5. The other three terms in $R_{12}$ can be similarly bounded. Thus, $R_{1}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$.
Symmetric arguments imply the analogous term for the control arm satisfies $R_{0}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$. Therefore, the results in Theorem 4 follow from the central limit theorem. $\square$